Find the focus of the parabola that has vertex (7, 1), opens horizontally, and passes through the point (11, 0).
Answer by lwsshak3(11628)

Find the focus of the parabola that has vertex (7, 1), opens horizontally, and passes through the point (11, 0)
**
Standard form of equation for given parabola:
(y-k)^2=4p(x-h), with (h,k) being the (x,y) coordinates of the vertex
..
For given equation:
(y-1)^2=4p(x-7)
using given point (11,0) to find 4p
4p=(0-1)/(11-7)=-1/4
p=-1/16
Equation: (y-1)^2=(-1/4)(x-7)
This is an equation of a parabola which opens leftwards, axis of symmetry, y=1
Ans: focus: (111/16,1) (p units from x-coordinate of vertex(7,1) on axis of symmetry)
note: 7=112/16